18. Solve
(rs)x-1=x-3
I have no idea on how to solve this problem and my head is beginning to hurt. Can someone help me please.

Answer by chitra(359)

The given expression is:

rs(x - 1) = x - 3

We remove the radical on the left hand side by squaring both sides of the given expression.

That is:

(x - 1) = (x - 3)^2

Expanding the right hand side by using the formula (a - b)^2, we get:

(x - 1) = x^2 - 6x + 9

Adding 1 - x to both sides of the above equation, we get:

x^2 - 6x + 9 + 1 - x = 0

x^2 - 7x + 10 = 0 [Adding the like terms, we get]

Using the method of Factorization, we get:

x^2 - 5x - 2x + 10 = 0

x(x - 5) - 2(x - 5) = 0

(x - 2)(x - 5) = 0

Thus, the values of x are 2 and 5.